Compute the determinant
4 0 -1
3 6 -2
-2 5 1

Answer by Boreal(15235)

Start writing out the first row, only the first two columns. Do the same for the second and third rows. This isn't required, but saves trying to remember what the third or the second and third numbers are. Go up to down right first and subtract from that the down to up and to the right second.
4 0 -1==4==0
3 6 -2==3==6
-2 5 1=-2===5
It is 4*6*1+0*-2*-2+-1*3*5=24+0-15=9
subtract the other way
-2*6*-1+5*-2*4+1*3*0=12-40+0=-28
Subtract 9-(-28)=37